OpenStudy (anonymous):

A certain region is losing 16 million acres of rain forest each year. Find a linear function f that calculates the change of acres in the rain forest in millions in x years. Find f(5) and interpret the result.

OpenStudy (australopithecus):

x(-1.6*10^7) = y

OpenStudy (australopithecus):

I made 1.6*10^7 negative because you are losing acreage

OpenStudy (australopithecus):

y = the total acreage of rainforest lost per year

OpenStudy (australopithecus):

so, f(x) = x(-1.6*10^7)

OpenStudy (australopithecus):

now solve f(5) by replacing x in the equation with 5

OpenStudy (australopithecus):

f(x) = x(-1.6*10^7) at f(0) = 0(-1.6*10^7) = 0 thus I dont see how your solution is any different from mine

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

@sauravshakya the only difference is, is I did not reduce the number to 16, which is probably the correct thing to do as it tells you to calculate in millions

OpenStudy (australopithecus):

so I guess we can agree that the solution is f(x) = x(-16) where the answer is in millions

OpenStudy (anonymous):

Okay Thanks let me see if I can get the same answer

OpenStudy (australopithecus):

0*-16 = 0 that is how

OpenStudy (australopithecus):

yes but its asking for the change in acres per year not the total acres

OpenStudy (australopithecus):

we know that it changes by -16 million per year, thus the change is just the function f(x) = -16x

OpenStudy (anonymous):

Oh yes..........

OpenStudy (australopithecus):

it is ok I'm guilty of not fully reading or comprehending a question often on this site :)

OpenStudy (anonymous):

I am having trouble with how much was lost per year. Is is 5 year and what rate?

OpenStudy (australopithecus):

f(5) just means you replace x with 5 in the equation

OpenStudy (australopithecus):

A point use to be (x,y) now it is (x,f(x)) in your case you are looking for the point (5,f(5))

OpenStudy (anonymous):

So it would be over a five years? Than you no point is needed. I am workin gon the rate per year. Would it be 80 million a year

OpenStudy (australopithecus):

f(x), g(x), s(x)... it is just new way of saying y, we use this notation for 2 reasons, it makes it easy to talk about multiple functions and 2 it allows us to talk about a specific point in a function easily that otherwise could be ugly in reality

OpenStudy (australopithecus):

-16*5 = f(5)

OpenStudy (anonymous):

Okay thank you

OpenStudy (australopithecus):

f(5) = -16*5 thus you would lose 80 million acres you have the point (5, -80)

OpenStudy (australopithecus):

if you have any further questions ask